# Conformal Prediction Adaptive to Unknown Subpopulation Shifts

### v1

@@ Abstract
Conformal prediction is widely used to equip black-box machine learning models with uncertainty quantification enjoying formal coverage guarantees. However, these guarantees typically break down in the presence of distribution shifts, where the data distribution at test time differs from the training (or calibration-time) distribution. In this work, we address subpopulation shifts, where the test environment exhibits an unknown and differing mixture of subpopulations compared to the calibration data. We propose new methods that provably adapt conformal prediction to such shifts, ensuring valid coverage without requiring explicit knowledge of subpopulation structure. Our algorithms scale to high-dimensional settings and perform effectively in realistic machine learning tasks. Extensive experiments on vision (with vision transformers) and language (with large language models) benchmarks demonstrate that our methods reliably maintain coverage and controls risk in scenarios where standard conformal prediction fails.
@@ 1 Introduction
In high-stakes real-world applications of machine learning, such as healthcare, uncertainty quantification (UQ) is crucial to safeguard patient health from the risks posed by model uncertainty. Conformal prediction techniques 10.5555/1062391 offer a framework for uncertainty quantification before model deployment. Conformal prediction operates similarly to a doctor’s differential diagnosis. A doctor gathers patient data and uses medical expertise to eliminate unlikely conditions, narrowing down to a list of probable diagnoses. Similarly, conformal prediction produces a set of possible labels, with a high probability that the true label lies within this set. The size of the prediction set reflects the model’s uncertainty—the larger the set, the greater the uncertainty, and vice versa.
Formally, conformal prediction guarantees marginal coverage, meaning that for a given input $X_{\text{test}}$ with unknown label $Y_{\text{test}}$ and a user-defined error rate $\alpha$, the probability that $Y_{\text{test}}$ lies in the prediction set $C_{\alpha}(X_{\text{test}})$ is at least $1-\alpha$, i.e.,
$Pr(Y_{\text{test}}\in C_{\alpha}(X_{\text{test}}))\geq 1-\alpha.$ (1)
The size of the prediction set $C_{\alpha}(X_{\text{test}})$ reflects the level of uncertainty–larger sets indicate higher uncertainty, while smaller sets signal greater confidence. The threshold used in conformal prediction determines how conservative the prediction set is, balancing between marginal coverage and uncertainty.
Standard conformal prediction offers provable marginal coverage guarantees under the assumption that test data is exchangeable with the training data. However, in many real-world scenarios, this assumption is violated due to distribution shifts. One of the most common types of distribution shift is subpopulation shift, where the proportions of subpopulations differ between training and deployment environments yang2023changehardcloserlook . A key challenge arises when different subpopulations present varying levels of prediction difficulty, requiring distinct thresholds to maintain reliable marginal coverage across all subpopulations. Distribution shifts, particularly subpopulation shifts, complicate this task further by causing the proportions of subpopulations to differ between training and test environments. As a result, a uniform threshold might not provide adequate marginal coverage for all subpopulations.
@@
  • We introduce a novel class of algorithms that are adaptive to arbitrary mixtures over domains by utilizing a learnt domain classifier. These can be seen as test-time adaptation methods that adaptively adjust the conformal prediction threshold. We prove that under mild assumptions, our new algorithms guarantee tight coverage for arbitrary subpopulation shifts.
  • We extend to the setting when we have no knowledge of the domains (either at calibration time or test time). In this case, we adaptively filter and reweight the calibration data to adaptively pick a threshold for each test data point.
  • We run extensive experiments simulating realistic subpopulation shifts on high-dimensional vision classification datasets. We show that our methods consistently provide tight coverage across test environments, whereas standard conformal prediction under or over-covers.
@@
Figure 1: (Left) Example of subpopulation shifts with 4 domains. Each colored square represents data from a particular domain. Train and test environments are mixtures of the same set of domains but at different proportion. Score distributions (gray for train environment and blue for each test environment) and threshold calculated from standard conformal prediction are shown for each train/test environment. Subpopulation shift for test environment 1 leads to roughly the ideal coverage while shifts for test environment 2 and 3 lead to significant under and over-coverage respectively. (Right) The same issue arises in LLM hallucination detection across different test environments. Standard LLM uncertainty estimation method (blue) is sensitive to distribution shifts displaying high variance in its hallucination detection recall across test-environments, while the recall with our modification (orange) tightly follows the desired target recall.
2 Problem Setup
2.1 Background
Notation.
We let $\mathcal{X}$ and $\mathcal{Y}$ denote the input and target space of a multiclass classification task. $\hat{f}:\mathcal{X}\rightarrow\Delta^{J}$ is the pre-trained classifier for the classification task and the output of $\hat{f}$ is a probability distribution over $J$ possible outcomes, e.g., the softmax output of a neural network. $\hat{f}(X)_{i}$ represents the $i$-th entry of the output of $\hat{f}$. We let $z:\mathcal{X}\rightarrow\mathbb{R}^{d}$ denote an embedding function that maps the input into a d-dimensional embedding space, and $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ stands for some similarity function between embeddings.
We denote by $\mathbb{P}_{k}$ the distribution of the $k$-th domain, where there are $K$ domains in total. The calibration dataset from the $k$-th domain is represented as $\{(X^{k}_{i},Y^{k}_{i});\hat{f}\}^{n_{k}}_{i=1}$, and each pair $(X^{k}_{i},Y^{k}_{i})$ is assumed to be drawn i.i.d. from $\mathbb{P}_{k}$. The overall calibration dataset is sampled from the training environment $\mathbb{P}_{\text{train}}$, which is a mixture of the $K$ domain distributions. The score function is represented by $S:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ and will make use of $\hat{f}$. We define $m_{k}(q)$ as |$|\{S(X^{k}_{i},Y^{k}_{i};\hat{f})\leq q\}|$, which is the number of number of calibration data in domain $k$ with score less than or equal to $q$.
Lastly, the test set is denoted by $\{(X^{i}_{\text{test}},Y^{i}_{\text{test}})\}_{i=1}^{n_{\text{test}}}$. These samples are drawn from a test environment $\mathbb{P}_{\text{test}}$, which is an unknown mixture of domains. We denote the set of all possible test environments by $\mathcal{D}$. Examples of such test environments are illustrated in Figure 1.
Score Functions
A score function maps an input pair $(X,Y)$ to a real-valued score. A larger score indicates less conformity between $(X,Y)$ and other training data. Conformal Prediction allows an arbitrarily defined score function, however, it could give uninformative prediction sets. For the vision task, we examine our methods with 3 score functions: Least Ambiguous Set-valued Classifier (LAC) Sadinle_2018 , Adaptive Prediction Set (APS) romano2020classificationvalidadaptivecoverage , and Regularized Adaptive Prediction Set (RAPS) angelopoulos2022uncertaintysetsimageclassifiers . For the language task, we examine 3 score functions: Length Normalized Scoring (LNS) malinin2021uncertainty , Meaning-Aware Response Scoring (MARS) bakman2024mars , and Degree Matrix Uncertainty lin2023generating .
Conformal Prediction Under Exchangeability
Given a test data $X_{\text{test}}$ with unknown label $Y_{\text{test}}$, a calibration set $\{(X_{i},Y_{i})\}^{n}_{i=1}$, which is distinct from train and test set, and a user defined error rate $\alpha$, the goal of conformal prediction is to build a prediction set $C_{\alpha}$ that satisfies 1. To conformalize a model to output a valid prediction set, the following procedure is followed: First, a score function $S$ is defined. Second, the threshold $\hat{q}_{\alpha}$ is computed as the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$ quantile of $\{S(X_{i},Y_{i};\hat{f})\}_{i=1}^{n}$. Lastly, the prediction set $C_{\alpha}(X_{test})$ is returned such that $C_{\alpha}(X_{test})=\{y|S(X_{test},y;\hat{f})\leq\hat{q}_{\alpha}\}$. If the calibration data and the test data are drawn i.i.d from the some domain, then $C(X_{test})$ satisfies the marginal coverage guarantee due to exchangeability between calibration and test data DBLP:journals/corr/abs-2107-07511 . We refer to this method as the standard or unweighted conformal prediction throughout the paper.
Conformal Risk Control
The conformal prediction framework can be extended to provide guarantee beyond coverage. Given a prediction set $C_{\hat{\lambda}}(X_{\text{test}})=\{y|S(X_{\text{test}},y)\leq\hat{\lambda}\}$, a loss function $\ell$ that decreases as $|C_{\hat{\lambda}}(X_{\text{test}})|$ increases, and a user defined error rate $\alpha$, the conformal risk control guarantee is defined as,
$\mathbb{E}[\ell(C_{\hat{\lambda}}(X_{\text{test}}),Y_{\text{test}})]\leq\alpha$ (2)
angelopoulos2023conformalriskcontrol . Note that the marginal coverage guarantee can be reduced to 2 if we define $\ell$ as the miscoverage loss, i.e., $\ell(C_{\hat{\lambda}}(X_{\text{test}}),Y_{\text{test}})=\mathbbm{1}\{Y_{\text{test}}\notin C_{\hat{\lambda}}(X_{\text{test}})\}$. We refer to angelopoulos2023conformalriskcontrol for the details of conformal risk control. One application for the conformal risk control framework is in large language model (LLM) uncertainty estimation, in particular, hallucination detection. Hallucination refers to when an LLM generate responses that are factually false or inconsistent with the training data. Conformal risk control can be used to select a threshold to determine whether an LLM output is a hallucination or not while maintaining a theoretical bound to metrics such as sensitivity or precision.
@@ 2.2 Conformal Prediction under Subpopulation Shifts
The standard conformal prediction procedure relies on the exchangeability assumption between calibration and test data. However, in many real-world scenarios—such as dynamic time series—this assumption often does not hold prinster2024conformalvalidityguaranteesexist . In this work, we focus on the setting of subpopulation shifts. Specifically, the calibration data is drawn from one of the $K$ in-domain distributions, denoted by $\mathbb{P}_{k}$. In contrast, the test data is sampled i.i.d. from a test environment $\mathbb{P}_{\text{test}}$, which is a mixture of these in-domain distributions, i.e.,
$\mathbb{P}_{\text{test}}=\sum^{K}_{k=1}\lambda_{k}\mathbb{P}_{k},$ (3)
@@
where $\lambda_{k}$ is the probability that $\mathbb{P}_{\text{test}}$ is drawn from $\mathbb{P}_{k}$. During test time, $\lambda_{k}$’s are unknown.
Failure of Standard Conformal Prediction.
@@
Under standard conformal prediction, the marginal coverage in Equation 1 is not guaranteed if the test data is not exchangeable with the calibration data due to subpopulation shifts. For instance, if the test environment has higher probability to be drawn from a harder domain, i.e., $\lambda_{k}$ is large for domain $k$ where data typically receive higher scores, then standard conformal prediction would result in under-coverage. Conversely, if $\lambda_{k}$ is large for domain $k$ which has data with lower scores, it would lead to over-coverage. As illustrated in Figure 1, test environment 2 exhibits under-coverage, while test environment 3 demonstrates over-coverage.
Prior Approaches
To tackle the issue caused by distribution shifts, we need to weigh calibration data from each domain accordingly based on the test environment. tibshirani2020conformalpredictioncovariateshift proposes a weighted version of conformal prediction for the case of generic covariate shifts, i.e., the input of the calibration data $X$ is drawn from some distribution $\mathcal{P}$ while the input of the test data $X_{n+1}$ is drawn from some distribution $\tilde{\mathcal{P}}$. The method weighs the score of each calibration data by its likelihood ratio, $d\tilde{\mathcal{P}}/d\mathcal{P}$ and is able to achieve 1. However, the method requires either (1) knowledge of $\tilde{\mathcal{P}}$ and $\mathcal{P}$ at test time or (2) an extra held out set drawn from the test environment to estimate $d\tilde{\mathcal{P}}/d\mathcal{P}$. For case 1, $\lambda$ is unknown at test time so $\tilde{\mathcal{P}}$ is not known while for case 2, test data might be scarce so an estimate of the likelihood ratio is not possible. Another method to adapt conformal prediction to distribution shifts is proposed by Cauchois_2024 . In that work, distribution shifts that make the score large are estimated and marginal coverage guarantee is provided for such shifts. To adapt this idea to our settings, the worst case shift is the case when the test environment is drawn from the hardest domain completely, i.e., the $\mathbb{P}_{\text{test}}=\mathbb{P}_{k}$ where the score from domain $k$ are typically larger than all other domains.
3 Adapting to subpopulation shifts with known domains
@@ 3.1 Weighted Conformal Prediction
To solve the issue caused by distribution shifts, we need to weigh calibration data from each domain differently based on the test environement. For example, if $\lambda_{k}$ is high, we will need to weight calibration data from domain $k$ higher since $\mathbb{P}_{k}$ represents the test environment more closely. To find the correct weighting, we propose Algorithm 1 which takes a separate model, $c:\mathcal{X}\rightarrow[0,1]^{K}$, named domain classifier, to predict the true $Pr((X_{\text{test}},Y_{\text{test}})\sim\mathbb{P}_{k}|X_{\text{test}})$ for each domain $k$. Theorem 3.1 states that if $c$ is a Bayes-optimal classifier, then we get marginal coverage guarantee.
@@
If $c$ is a Bayes-optimal classifier, the output of Algorithm 1 $C_{\alpha}$, satisfies $(1-\alpha)$ coverage as in eqn. (1).
The proof of Theorem 3.1 can be found in Appendix A.1. Intuitively, if we have a Bayes-optimal domain classifier, the weight given to the domains which are more likely should be higher. In fact, by weighting the calibration scores based on $\lambda$, we can adopt the partial exchangeability proof of lu2023federatedconformalpredictorsdistributed to prove our claim. In the extreme case where the test environment is one of the $K$ in-distribution domains, i.e. $(X_{\text{test}},Y_{\text{test}})\sim\mathbb{P}_{\text{test}}=\mathbb{P}_{k}$, we have that $\hat{\lambda}(X_{\text{test}})_{i}=1$ for $i=k$ and 00 otherwise. We see that the Algorithm 1 reduces to the case of standard conformal prediction which satisfies 1 since $(X_{\text{test}},Y_{\text{test}})$ is now exchangeable with calibration data from domain $k$.

### v2

@@ Abstract
Conformal prediction is widely used to equip black-box machine learning models with uncertainty quantification, offering formal coverage guarantees under exchangeable data. However, these guarantees fail when faced with subpopulation shifts, where the test environment contains a different mix of subpopulations than the calibration data. In this work, we focus on unknown subpopulation shifts where we are not given group-information i.e. the subpopulation labels of datapoints have to be inferred. We propose new methods that provably adapt conformal prediction to such shifts, ensuring valid coverage without explicit knowledge of subpopulation structure. While existing methods in similar setups assume perfect subpopulation labels, our framework explicitly relaxes this requirement and characterizes conditions where formal coverage guarantees remain feasible. Further, our algorithms scale to high-dimensional settings and remain practical in realistic machine learning tasks. Extensive experiments on vision (with vision transformers) and language (with large language models) benchmarks demonstrate that our methods reliably maintain coverage and effectively control risks in scenarios where standard conformal prediction fails.
@@ 1 Introduction
In high-stakes real-world applications of machine learning, such as healthcare, uncertainty quantification (UQ) is crucial to safeguard patient health from the risks posed by model uncertainty. Conformal prediction (CP) techniques (Vovk et al., 2005) offer a framework for uncertainty quantification before model deployment. Formally, conformal prediction guarantees marginal coverage, meaning that for a given input $X_{\text{test}}$ with unknown label $Y_{\text{test}}$ and a user-defined error rate $\alpha$, the probability that $Y_{\text{test}}$ lies in the prediction set $C_{\alpha}(X_{\text{test}})$ is at least $1-\alpha$, i.e.,
$Pr(Y_{\text{test}}\in C_{\alpha}(X_{\text{test}}))\geq 1-\alpha,\ \text{ for }(X_{\text{test}},Y_{\text{test}})\sim\mathbb{P}_{\text{test}}\,.$ (1)
The size of the prediction set $C_{\alpha}(X_{\text{test}})$ reflects the level of uncertainty–larger sets indicate higher uncertainty, while smaller sets signal greater confidence. The threshold used in conformal prediction determines how conservative the prediction set is, balancing between coverage and uncertainty.
Standard conformal prediction offers provable marginal coverage guarantees under the assumption that test data is exchangeable with the training data. However, in many real-world scenarios, this assumption is violated due to distribution shifts. One of the most common types of distribution shift is subpopulation shift, where the proportions of subpopulations differ between training and deployment environments (Yang et al., 2023). A key challenge arises when different subpopulations present varying levels of prediction difficulty, requiring distinct thresholds to maintain reliable marginal coverage across all subpopulations. Distribution shifts, particularly subpopulation shifts, complicate this task further by causing the proportions of subpopulations to differ between training and test environments. As a result, a uniform threshold might not provide adequate marginal coverage for all subpopulations.
@@
  • We introduce an algorithm class that is adaptive to arbitrary mixtures over domains by utilizing a learned domain classifier. These can be seen as test-time adaptation methods that adaptively adjust the conformal prediction threshold. We prove that under mild assumptions, our new algorithms guarantee tight coverage for arbitrary subpopulation shifts.
  • We extend the method to when we do not have access to (even imperfect) domain classifier. In this case, we adaptively filter and reweight the calibration data to adaptively pick a threshold for each test data point.
  • We run extensive experiments simulating realistic subpopulation shifts on high-dimensional vision classification datasets. We show that our methods consistently provide tight coverage across test environments, unlike prior approaches that under or over cover.
@@
Figure 1: (Left) Example of subpopulation shifts with 4 domains and 3 test environments. Each colored square represents data from a particular domain. Train and test environments are mixtures of the same set of domains but at different proportion. Score distributions (gray for train environment and blue for each test environment) and threshold calculated from standard conformal prediction are shown for each train/test environment. Subpopulation shifts leads to roughly the ideal coverage in test environment 1, whereas shifts for test environment 2 and 3 lead to significant under and over-coverage respectively. (Right) The same issue arises in LLM hallucination detection across different test environments. Standard LLM uncertainty estimation method (blue) is sensitive to distribution shifts displaying high variance in its hallucination detection recall, while the recall with our modification (orange) tightly follows the desired target recall.
2 Limitations of Prior Approaches for Subpopulation Shifts
@@ 2.1 Conformal Prediction under Subpopulation Shifts
The standard conformal prediction procedure relies on the exchangeability assumption between calibration and test data. However, in many real-world scenarios—such as dynamic time series—this assumption often does not hold (Prinster et al., 2024). In this work, we focus on the setting of subpopulation shifts. Specifically, we have $K$ domain-specific distributions, denoted by $\mathbb{P}_{k}$. The test data is sampled i.i.d. from a test environment $\mathbb{P}_{\text{test}}$ i.e.,
$\mathbb{P}_{\text{test}}=\sum^{K}_{k=1}\lambda_{k}\mathbb{P}_{k},\vskip-5.0pt$ (2)
@@
where $\lambda_{k}$ is the probability that $\mathbb{P}_{\text{test}}$ is drawn from $\mathbb{P}_{k}$. Importantly, the weights $\lambda_{k}$’s are unknown and arbitarily different from the mixture weights of the calibration data distribution.
@@
Failure of Standard Conformal Prediction (CP). In standard CP, (1) is not guaranteed if the test data is not exchangeable with the calibration data due to subpopulation shifts (Tibshirani et al., 2020). For instance, if the test environment has higher probability to be drawn from a harder domain, i.e., $\lambda_{k}$ is large for domain $k$ where data typically receive higher scores, then standard conformal prediction would result in under-coverage. Conversely, if $\lambda_{k}$ is large for domain $k$ which has data with lower scores, it would lead to over-coverage. As illustrated in Figure 1, test environment 2 exhibits under-coverage, while test environment 3 demonstrates over-coverage. See App. B for an overview of notation and background on CP.
CP under Distribution Shifts. When the distribution shift is known, Tibshirani et al. (2020) showed that we can recover coverage by reweighting calibration score by the covariate likelihood ratio between training and test distribution. However, this approach relies on either knowledge of the test covariate distribution - which in our case is unknown - or estimating this density ratio using a held-out set sampled from the test distribution, which is prohibitive in high-dimensional modern ML. Alternatively, robust or max CP proposes to use a fixed threshold that guarantees coverage for a worst-case distribution shift (Cauchois et al., 2024). In our setup, this corresponds to $\mathbb{P}_{\text{test}}$ being drawn only from the ”hardest” domain. Suppose $\hat{q}_{\alpha}^{k}$ is the domain-specific threshold for $\mathbb{P}_{k}$ i.e. $\hat{q}_{\alpha}^{k}=\lceil(n_{k}+1)(1-\alpha)/n_{k}\rceil$-quantile of the $n_{k}$ calibration data from domain $k$. The max method returns prediction set using the score threshold $\hat{q}_{\alpha}:=\max_{k\in[K]}\hat{q}^{k}_{\alpha}.$ However, as we later show, this can be conservative and have significant over-coverage. We want to be adaptive to the actual difficultly of $\mathbb{P}_{k}$ instead.
Group Conditional CP. Group-conditional CP is a closely related approach that has received significant recent attention (Jung et al., 2022; Gibbs et al., 2024; Kiyani et al., 2024; Bairaktari et al., 2025) ,which states that given the input space $\mathcal{X}$ and a collection of groups $\mathcal{G}\subseteq 2^{\mathcal{X}}$, for all $G\in\mathcal{G}$,
$Pr(Y_{\text{test}}\in C_{\alpha}(X_{\text{test}})|X_{\text{test}}\in G)\geq 1-\alpha.$ (3)
This strengthens the standard marginal CP guarantee and simultaneously provides coverage for a collection of subset of $\mathcal{X}$, and is itself a tractable relaxation of exact conditional coverage which is known to be impossible (Barber et al., 2020). Suppose we can satisfy equation 3 for each of the $k$ domains with $\mathcal{G}=[K]$ i.e. for any $G=\mathbb{P}_{k}$. Such as $C_{\alpha}$ would also satisfy coverage for our subpopulation shift setting for any weight vector $\lambda$ in equation 2 and hence solves our problem. However, as we next argue, this approach has serious limitations.
2.2 Limitations of Conditional Conformal Prediction
Conditional CP needs group membership. Indeed, satisfying group-conditional coverage implies coverage in our settings, if we define the collection of groups to be the collection of the $K$ domains. However, at test time, group-conditional CP critically needs knowledge of group membership, which is rarely available in practice. Methods have been proposed to learn group memberships with predefined groups (Gibbs et al. (2024), Jung et al. (2022)) or learn the natural partition of the input space (Kiyani et al., 2024), but there has not yet been an extensive theoretical or practical investigation of what to do when group membership information is imperfect during test time. Similar to our proposed methods, Gibbs et al. (2024) also employs a two-stage approach where they also train a domain classifier. However, their analysis still assumes perfect group information and leaves open the effect of an imperfect domain classifier.
Conditional CP coverage degrades with imperfect group membership.
Theorem 2.1.
Suppose we are given a algorithm $C_{\alpha}$ that given perfect group information (whether $X_{\text{test}}\sim\mathbb{P}_{k}$) obtains perfect domain-conditional coverage as in equation 3. Then, there exist domain distributions $\{\mathbb{P}_{k}\}_{k\in[K]}$, and a domain classifier $c(X_{\text{test}})\in[K]$ with conditional accuracy $\gamma\in[0,1]$ for every domain $k$, such that if we use $c(X_{\text{test}})$ as our imperfect group-information, then
$Pr(Y_{\text{test}}\in C_{\alpha}(X_{\text{test}})|X_{\text{test}}\sim\mathbb{P}_{k})\leq\max(0,\gamma-\alpha))\,.$
The above theorem shows that the coverage guarantees of any group-conditional conformal predictor can significantly degrade when paired with imperfect group information, demonstrating a big drawback relying in realistic settings where such information is unlikely to be given. This limitation is not just theoretical as we show on the right. We evaluate our methods against the two-stage group-conditional approach (Gibbs et al., 2024) (referred to as Conditional Calibration (CC)) on 100 test environments with varying distribution shift and show that it has significantly higher mis-coverage compared with our methods (A1, A2, and oracle), and is even worse than standard unweighted CP for small $1-\alpha$!
3 Subpopulation shifts with an imperfect domain classifier
@@ 3.1 Weighted Conformal Prediction
To solve the issue caused by distribution shifts, we need to weigh calibration data from each domain differently based on the test environment. For example, if $\lambda_{k}$ is high, we will need to weigh calibration data from domain $k$ higher since $\mathbb{P}_{k}$ represents the test environment more closely. To find the correct weighting, we propose Algorithm 1 which takes a separate model, $c:\mathcal{X}\rightarrow\Delta^{K}$, named domain classifier, to predict the true $Pr((X_{\text{test}},Y_{\text{test}})\sim\mathbb{P}_{k}|X_{\text{test}})$ for each domain $k$. Theorem 3.1 states that if $c$ is a Bayes-optimal classifier, then we get marginal coverage guarantee.
@@
Suppose $c:\mathcal{X}\rightarrow\Delta^{K}$ is a domain classifier that maps the input to a probability distribution over the $K$ domains and $(X_{\text{test}},Y_{\text{test}})$ is sampled from $\mathbb{P}_{\text{test}}$, as defined in section 2.1. If $c$ is a Bayes-optimal classifier, the output of Algorithm 1, $C_{\alpha}$, satisfies
$Pr(Y_{\text{test}}\in C_{\alpha}(X_{\text{test}}))\geq 1-\alpha.$
We refer to Algorithm 1 using a Bayes-optimal domain classifier as the oracle method.
The proof of Theorem 3.1 can be found in Appendix A.2. Intuitively, if we have a Bayes-optimal domain classifier, the weight given to the domains which are more likely should be higher. In fact, by weighting the calibration scores based on $\lambda$, we can adopt the partial exchangeability proof of Lu et al. (2023) to prove our claim. In the extreme case where the test environment is one of the $K$ in-distribution domains, i.e. $(X_{\text{test}},Y_{\text{test}})\sim\mathbb{P}_{\text{test}}=\mathbb{P}_{k}$, we have that $\hat{\lambda}(X_{\text{test}})_{i}=1$ for $i=k$ and 0 otherwise. We see that the Algorithm 1 reduces to the case of standard conformal prediction which satisfies (1) since $(X_{\text{test}},Y_{\text{test}})$ is now exchangeable with calibration data from domain $k$.